OpenStudy (anonymous):

find the differnece [-10]-[-3]? and how did u do it?

OpenStudy (anonymous):

The steps are (-10)-(-3) =-10+3 =-7

OpenStudy (anonymous):

Are you meaning [-10] as absolute value?

OpenStudy (anonymous):

no parenthesis

OpenStudy (anonymous):

oh, then yeah. Just do the multiplication of the - on the outside to the -3, then sum.

OpenStudy (anonymous):

Hey Polpak, how old are you?

OpenStudy (anonymous):

Old enough. (I hope)

OpenStudy (anonymous):

I just want to know. That's it. Could you please tell.

OpenStudy (anonymous):

I think I'm 33.

OpenStudy (anonymous):

\(33 \pm 1\)

OpenStudy (anonymous):

Oh, I see. As everyone in this group was around 12-15, I thought so was the moderator of the group. Anyway, it's nice to know that it's not being managed by a kid.

OpenStudy (anonymous):

What is the uncertainty for?

OpenStudy (anonymous):

Ah, no. There are a few of us old people around here.